OpenStudy (anonymous):

Hi everyone, I need help with the following problem: In a particular city, the monthly supply and demand for a new type of tennis racket pf price p ( in $) and q (quantity) are modeled by the equatations: {S(p)= 38p - 1710 { D(q)= -46p + 4002 $ 45less or equal p less or equal $87. What is the quilibrium point? Thanks :)

OpenStudy (phi):

does D(q)= -46 q +4002 or -46 p + 4002

OpenStudy (anonymous):

Oops, sorry, it is D(p)= -46+4002

OpenStudy (phi):

you mean D(p)= -46p +4002 ? if so, set the 2 equations equal, and solve for p

OpenStudy (anonymous):

Please explain, I really don't get this at all...

OpenStudy (phi):

S(p) goes up a p goes up. D(p) goes down as p goes up. the meet somewheres. 

OpenStudy (anonymous):

something like this 38p -1710=-46p+4002 38p+46p=1710+4002 84p=5712 p=68

OpenStudy (phi):

the only thing to add, is that D(q)= -46p + 4002 $ 45less or equal p less or equal $87. so the equation is only valid for $45 <= p <= $87 so you should note that p is between these values.

OpenStudy (anonymous):

Thank for your help :)